the length of a rectangle is 4 centimeters less than triple the width and the area is 32 find the dimensions.
Answer by flame8855(424)

l= 3w-4
lw=32
(3w-4)(w) = 32
3w^2 - 4w -32 =0
(w-4)(3w+8) = 0
w = 4
lw=32 cm
l = 8 cm